# Hierarchical Large Scale Multirobot Path (Re)Planning

### v1

@@ VI Multi-commodity Flow with Optimal Detour
Our hierarchical approach relies on high-level planning to 1) regulate cell congestion and 2) preserve the bounded-suboptimality of inter-cell routing solutions. Thus, our high-level planner simplifies MAPF instances within cells and leads to real-time replanning. We abstract the partition as a directed graph $\mathcal{G}_{p}\!\!=\!\!\left(V_{p},E_{p}\right)$, where the vertices (nodes) represent cells and edges connect neighboring cells that share at least one face. Edges are weighted according to Euclidean distance between the cells’ centers of mass. The high-level planner finds a routing $\mathcal{U}^{i}\!\!=\!\!\left[P^{i}_{s},\cdots,P^{i}_{g}\right]$ for each robot $r^{i}$, where $P^{i}_{s}$ and $P^{i}_{g}$ are its start and goal cell, satisfying: 1) the influx of cell $m$ is under a user-defined value $\theta_{m}$, and 2) $cost(\mathcal{U}^{i})\!\!\leq\!\!w_{mcf}\cdot cost(\mathcal{U}^{i,*})$. Here, $w_{mcf}\!\!\geq\!\!1$ is a scalar representing the suboptimality bound for the routing solutions. $g^{i,*}$ is the optimal inter-cell routing of robot $r^{i}$. The influx of a node represents the number of robots entering the cell; we define influx formally in the following section.
@@ VI-A High-level Planning Formulation
The SOTA partition-based MAPF solvers [11, 12] suffer from cell congestion, leading to hard MAPF instances in certain cells, causing computational bottlenecks. To address this, we formulate the inter-cell routing as a variant of the MCF problem and propose multi-commodity flow with optimal detour (MCF/OD), which optimally distributes robots among cells such that the number of robots entering any intermediate cell $m$ (not the start or goal cells of the robot teams) is under a user-defined value $\theta_{m}$, if solution exists.
Specifically, robots sharing the same start cell $P_{s}$ and goal cell $P_{g}$ are one commodity $c_{sg}$. The commodities set $C\!\!=\!\!\left\{c_{1},\cdots,c_{O}\right\}$ includes all commodities given the robot positions. Solving the MCF problem results in optimal flows $\{y^{*}_{sgml}\}$, that is the number of robots in commodity $c_{sg}\!\!\in\!\!C$ traversing along edge $e_{ml}\!\in\!E_{p}$. In our minimal influx MCF formulation, the optimal flow solutions lead to minimized intermediate cell influx and the most dispersed routing. The cell influx of $P_{l}$ is defined as the total number of robots entering this cell, that is, $\sum_{c_{sg},e_{ml}}y_{sgml}$. By constraining the flows $y_{sgml}$ to the shortest paths between the start and goal cells $SP_{sg}$, as in (1e), (1f), we guarantee solution optimality. We formulate the minimal influx MCF problem as the following integer linear program (ILP):
@@
$y_{sgml}$ $=0,~{}\forall e_{ml}\notin SP_{sg},\forall c_{sg}\in C\vspace{-0.5em}$ (1f)
where $\mathcal{L}_{sg}$ represents the maximum flow for commodity $c_{sg}\in C$, and $\mathcal{L}_{in}$ represents the maximum influx among all the cells. By minimizing both, the objective function penalizes the maximum congestion among all cells and disperses the flows related to one commodity. We weight $\mathcal{L}_{sg}$ and $\mathcal{L}_{in}$ with coefficients $\alpha$ and $\beta$, respectively, and set $\beta\gg\alpha$ to prioritize minimizing $\mathcal{L}_{in}$. In (1d), $\mathbf{I}_{v^{i}}=\left[I_{y_{1}},\cdots,I_{y_{F}}\right]$, where $y_{f}\in\left\{y_{sgml}\right\},I_{y_{f}}\in\left\{0,1\right\}$ is an indicator function that returns $1$ when $y_{f}$ has positive flow to an intermediate vertex $v^{i}\in V_{p}$. $\mathbf{y}$ is the vector of all flows.
Despite the minimized influx to the intermediate cells, the minimal influx MCF formulation leads to congestion in certain cells due to tight constraints on the shortest paths (robots’ shortest paths may intersect at certain cells). To detour the robots optimally, ensuring that the number of robots entering any cell $m$ remains below its influx limit $\theta_{m}$, we present a complete and optimal solver, MCF/OD, in Algo. 1. It maintains a conflict tree and resolves congestion iteratively. The function $congestionDetection(\mathcal{G}_{p},P.solution,\boldsymbol{\theta})$ computes the influx for each cell in $\mathcal{G}_{p}$, given the flow solution $P.solution$ and returns the set of cells with their influx larger than the corresponding limit. The function $getAllConflict(P)$ returns the set of all commodities passing congested cells.
@@ VII Low-level Planner
Within each cell, the low-level planner, or the cell planner, computes collision-free paths that navigate robots to their local goals in an anytime fashion. The cell planner can be divided into three steps: 1) local goal assignment to determine the goal for each robot in a cell, 2) anytime MAPF/C to generate discrete paths, and 3) cell-crossing protocol for non-stop transiting between cells.
VII-A Local Goal Selection
At the beginning of cell planning, local goal selection aims to assign robots to the closest local goals while spreading out robots in an optimal manner by solving the following ILP:
@@ VII-A Local Goal Selection
where $A_{ij}\!\!\in\!\!\{0,1\}$ indicates if robot $r^{i}$ is assigned to local goal $lg^{j}$, and $D_{ij}$ is the Euclidean distance between $r^{i}$ and $lg^{j}$. Auxiliary variables $u_{j}$ in the objective function minimize the number of robots queueing at local goal $lg^{j}$, prioritizing filling less congested local goals first. Auxiliary variable $U$ in the objective function minimizes the maximum number of robots waiting in queue among all the local goals. This leads to evenly routing robots to different local goals to reduce congestion. A local goal is occupied if assigned with at least one robot. Thus, the number of robots waiting in queue for a local goal $lg^{j}$ is $\left(\sum_{i}A_{ij}-1\right)$.
@@ VII-B Anytime MAPF/C
We adopt anytime MAPF, which iteratively improves solution quality until a solution is needed, to facilitate real-time replanning. We use the current SOTA anytime MAPF LNS [20], which iteratively replans for a subset of robots and adapts to an improved solution if found. We use ECBS [21] as the initial planner as it provides a bounded-suboptimal solution and prioritized planning with SIPP [22] to rapidly iterate plans. We extend both ECBS and SIPP using MAPF/C to account for generalized conflicts.
@@ VII-B1 SIPP with generalized conflicts
SIPP compresses the time dimension into sparse safety intervals to significantly reduce the search space. The SIPP configuration augments the position with its safety intervals. In the resulting configuration space, $A^{*}$ finds the shortest path for a robot. Planned robots are considered moving obstacles and modify the safety interval of the traversed states. We propose SIPP with generalized conflicts (SIPP/C) with the following modifications and a different getSuccessors(s) algorithm, where the highlighted part differs from the original algorithm. In SIPP/C, the collision intervals, the complements of safety intervals, are added to vertices and edges based on the following vertex-vertex, edge-edge, and vertex-edge conflicts:
SIPP/C vertex-vertex conflict: for a robot at the vertex $u_{k}$ in a planned path at time $k$, we add the collision interval $[k,k]$ to vertices and edges that intersect with the robot-robot collision model at $pos(u_{k})$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$, $\forall v,e$. Note here, we use the swept model for $\mathcal{R}^{*}_{\mathcal{R}}(e)$ in SIPP.
SIPP/C edge-edge conflict: for a robot traversing an edge $e_{k}\!\!:=\!\!(u_{k},u_{k+1})$ at time $k$, we add the collision interval $[k,k]$ to vertices and edges that intersect with the robot’s swept collision model along edge $e_{k}$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}^{*}_{\mathcal{R}}(e_{k})$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}^{*}_{\mathcal{R}}(e_{k})$, $\forall v,e$.
SIPP/C edge-vertex conflict: for a robot to be stationary at the vertex $u_{k}\!\!=\!\!u_{k+1}$, we add the collision interval $[k,k]$ to vertices and edges that intersect with the robot’s swept collision model along edge $e_{k}\!\!:=\!\!(u_{k},u_{k+1})$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$, $\forall v,e$. Note that this degenerates to vertex-vertex conflict.
@@ 16             $s^{{}^{\prime}}$ = state of configuration $cfg$ with interval $i$ and time $t$
In Algo. 2, $E(s)$ is the action space at state $s$. Given the discrete path of robot $r^{i}$, we assign a time $t_{k}=k\Delta t$ to each discrete timestep and obtain the path $f^{i}$. $\Delta t$ is a user-defined value to satisfy the robot’s dynamic constraints. The low-level replanning happens every $\delta_{l}$ time interval.
@@ VII-C Cell-crossing Protocol
A robot idles at a local goal if the path in the next cell is not yet computed. To allow non-stop execution, we propose a cell-crossing protocol that results in robots always having a planned path to execute, even when traversing between cells. We buffer the hyperplane $H_{ml}$ by a distance $d_{e}$ towards cell $P_{m}$. All robots within the buffer will compute paths for their next cell before leaving the current one. Buffering is achieved by changing the hyperplane offset to $\mathcal{H}^{{}^{\prime}}_{a}=\mathcal{H}_{a}-\mathcal{H}_{\mathbf{n}}\cdot d_{e}$. By enforcing the buffer distance $d_{e}\geq\delta_{l}\cdot V_{max}$, where $V_{max}$ is the maximum speed of the robot, the robot is guaranteed to have a plan computed at least once before leaving its current cell. A robot entering the buffer zone will then have a plan to exit its current cell and transition through its next cell. The robot concurrently computes plans for its current and next cell, then concatenates them to form a complete transition plan. The robot then fixes this plan to lock the local goal and expected arrival time to next cell. Thus, when computing the plan for the next cell, the robot’s expected start time and position will be pre-determined and independent of cell planning order. Fig. 5 depicts our cell-crossing protocol.
@@ VIII Results and Discussion
Fig. 6 depicts typical solutions of the proposed algorithm. We summarize quantitative results in Table I, where MCF refers to MCF/OD and One-shot MCF running in parallel, as described in Sec. VI-A. Note that, as $\left|V\right|$ and $\left|E\right|$ suggest, the number of vertices and edges increases after partitioning as we add local goals and corresponding edges. With a small computational overhead $\bar{t}_{\mathrm{high}}$, the proposed high-level planner effectively reduces the congestion among cells compared to both greedy and partitionless baseline approaches, by inspecting $\bar{N}_{\mathrm{max}}$. Here $\bar{t}_{\mathrm{high}}$ is the average high-level planning time and $\bar{N}_{\mathrm{max}}$ is the averaged maximum number of robots in a cell throughout the whole execution. By increasing the number of cells $Q$, the algorithm significantly reduces MAPF computation time. Specifically, for the average low-level planning time $\bar{t}_{\mathrm{low}}$, in instance “Circle74”, the proposed algorithm runs $616$-times faster than the baseline method and $544$-times faster in instance “Circle142”. Both the average low-level replanning time $\bar{t}_{\mathrm{low}}$ and the averaged maximum low-level replanning time $\bar{t}^{\mathrm{max}}_{\mathrm{low}}$ are within the real-time regime for all instances. Since we use an anytime algorithm, we only record its initial planning time in the Table, which can be directly compared to the baseline. As we would expect, while performing in real-time, the algorithm yields suboptimal solutions compared to the baseline MAPF solution, according to the average makespan $\bar{T}$. Because the partitioning invalidates the global optimality of MAPF algorithms, and the high-level planner detours robots among the partition, lengthening planned paths.
@@ VIII-A Effectiveness of Partition in Low-level Planning
Fig. 7a shows a quantitative evaluation of the low-level planning time and its solution makespan when changing the number of cells in “Circle74”. As a robust statistic, we measure the median low-level planning time. As the result suggests, the logarithm of low-level planning time decreases significantly as we increase the number of cells. The result indicates that by partitioning the workspace and parallelizing computation, the low-level planning time decreases significantly. Furthermore, the makespan increases at the beginning then plateaus as the number of cells increases. This is expected since as the number of cells increases, first the high-level planner is more effective in detouring the robots to reduce congestion. At a certain point, no more detour is required since congestion is regulated. Additionally, the partition breaks the global optimality of the MAPF planner, leading to degeneracy in solution quality.
@@
Our algorithm runs MAPF in parallel for all cells; in Fig. 7b, we investigate the utilization of multi-threading. The orange dashed curve represents the theoretical lower bound of the total low-level planning time (assuming no overhead is introduced when allocating the computation to different threads). We observe the algorithm computation time generally follows the same curve, while the gap between the theoretical lower bound and experimental computation time increases as we increase the number of threads, indicating an increase in overhead for parallel computation in more cells.

### v2

@@ VI Multi-commodity Flow with Optimal Detour
Our hierarchical approach relies on high-level planning to 1) regulate cell congestion and 2) preserve the bounded-suboptimality of inter-cell routing solutions. Thus, our high-level planner simplifies MAPF instances within cells and leads to real-time replanning. We abstract the partition as a directed graph $\mathcal{G}_{p}\!\!=\!\!\left(V_{p},E_{p}\right)$, where the vertices (nodes) represent cells and edges connect neighboring cells that share at least one face. Edges are weighted according to Euclidean distance between the cells’ centers of mass. The high-level planner finds an inter-cell routing $\mathcal{U}^{i}\!\!=\!\!\left[P^{i}_{s},\cdots,P^{i}_{g}\right]$ for each robot $r^{i}$, where $P^{i}_{s}$ and $P^{i}_{g}$ are its start and goal cell, satisfying: 1) the influx of cell $m$ is under a user-defined value $\theta_{m}$, and 2) $cost(\mathcal{U}^{i})\!\!\leq\!\!w_{mcf}\cdot cost(\mathcal{U}^{i,*})$. Here, $w_{mcf}\!\!\geq\!\!1$ is a scalar representing the suboptimality bound for the routing solutions. $\mathcal{U}^{i,*}$ is the optimal routing of robot $r^{i}$. The influx of a node represents the number of robots entering the cell; we define influx formally in the following section.
@@ VI-A High-level Planning Formulation
The SOTA partition-based MAPF solvers [13, 14] suffer from cell congestion, leading to hard instances in certain cells, causing computational bottlenecks. To address this, we formulate the inter-cell routing as a variant of the MCF problem and propose multi-commodity flow with optimal detour (MCF/OD), which optimally distributes robots among cells such that the number of robots entering any intermediate cell $m$ (not the start or goal cells of the robot teams) is under a user-defined value $\theta_{m}$, if a solution exists.
Specifically, robots sharing the same start cell $P_{s}$ and goal cell $P_{g}$ are one commodity $c_{sg}$. The commodities set $C\!\!=\!\!\left\{c_{1},\cdots,c_{O}\right\}$ includes all commodities given the robot positions. Solving the MCF problem results in optimal flows $\{y^{*}_{sgml}\}$, that is the number of robots in commodity $c_{sg}\!\!\in\!\!C$ traversing along edge $e_{ml}\!\in\!E_{p}$. In our minimal influx MCF formulation, the optimal flow solutions lead to minimized intermediate cell influx and the most dispersed routing. We define the cell influx of $P_{l}$ as the total number of entering robots, that is, $\sum_{c_{sg}\in C,e_{ml}\in E_{p}}y_{sgml}$. We formulate the following integer linear program (ILP):
@@
$y_{sgml}$ $=0,~{}\forall e_{ml}\notin SP_{sg},\forall c_{sg}\in C.$ (1f)
Here, $\mathcal{L}_{sg}$ represents the maximum flow for commodity $c_{sg}\in C$, defined by (1c). $\mathcal{L}_{in}$ represents the maximum influx among all the cells, defined by (1d), where $\mathbf{I}_{v^{i}}=\left[I_{y_{1}},\cdots,I_{y_{F}}\right]$, where $y_{f}\in\left\{y_{sgml}\right\},I_{y_{f}}\in\left\{0,1\right\}$ is an indicator function that returns $1$ when $y_{f}$ has positive flow to an intermediate vertex $v^{i}\in V_{p}$. $\mathbf{y}$ is the vector of all flows. By minimizing both, the objective function penalizes the maximum congestion among all cells and disperses the flows related to one commodity. We weight $\mathcal{L}_{sg}$ and $\mathcal{L}_{in}$ with coefficients $\alpha$ and $\beta$, respectively, and set $\beta\gg\alpha$ to prioritize minimizing $\mathcal{L}_{in}$. (1b) is the set of constraints for MCF formulation. The set of constraints (1e, 1f) enforces the flows $y_{sgml}$ onto the shortest paths between the start and goal cells $SP_{sg}$ guaranteeing the solution optimality.
Despite the minimized influx to the intermediate cells, the minimal influx MCF formulation leads to congestion in certain cells due to tight constraints on the shortest paths (robots’ shortest paths may intersect at certain cells). To detour the robots optimally, ensuring that the number of robots that enter any intermediate cell $m$ remains below its influx limit $\theta_{m}$ and no unnecessary detour is introduced, we present a complete and optimal solver, MCF/OD, in Algo. 1. It maintains a conflict tree and resolves congestion iteratively. The function $congestionDetection(\mathcal{G}_{p},P.solution,\boldsymbol{\theta})$ computes the influx for each cell in $\mathcal{G}_{p}$, given the flow solution $P.solution$ and returns the set of cells with their influx larger than the corresponding limit. The function $getAllConflict(P)$ returns the set of all commodities passing congested cells.
@@ VII Low-level Planner
Within each cell, the low-level planner, or the cell planner, computes collision-free paths that navigate robots to their local goals in an anytime fashion. The cell planner can be divided into three steps: 1) local goal assignment, 2) anytime MAPF/C generates discrete paths, and 3) cell-crossing protocol for non-stop transiting between cells.
VII-A Local Goal Assignment
As the first step, local goal assignment aims to route robots to the closest local goals while spreading out robots optimally by solving the following ILP:
@@ VII-A Local Goal Assignment
where $A_{ij}\!\!\in\!\!\{0,1\}$ indicates if robot $r^{i}$ is assigned to the $j$-th local goal, denoted as $lg^{j}$. $D_{ij}$ is the Euclidean distance between $r^{i}$ and $lg^{j}$. Auxiliary variables $u_{j}$ in the objective function minimize the number of robots queueing at local goal $lg^{j}$, prioritizing filling less congested local goals first. Auxiliary variable $U$ in the objective function minimizes the maximum number of robots waiting in queue among all the local goals. This leads to evenly routing robots to different local goals to reduce congestion. A local goal is occupied if assigned with at least one robot. Thus, the number of robots waiting in queue for a local goal $lg^{j}$ is $\left(\sum_{i}A_{ij}-1\right)$.
@@ VII-B Anytime MAPF/C
We adopt the SOTA anytime MAPF method, namely LNS [22], which iteratively improves the solution quality until a solution is needed, to facilitate real-time replanning. We use ECBS [23] as the initial planner as it provides a bounded-suboptimal solution and prioritized planning with SIPP [24] to rapidly replan for a subset of robots. We extend both ECBS and SIPP using MAPF/C to account for the robot embodiment.
@@ VII-B1 SIPP with generalized conflicts
SIPP compresses the time dimension into sparse safety intervals to significantly reduce the search space. The SIPP configuration augments the position with its safety intervals. In the resulting configuration space, $A^{*}$ finds the shortest path for a robot. Planned robots are considered moving obstacles and modify the safety interval of the traversed states. We propose SIPP with generalized conflicts (SIPP/C) with the following modifications and a different getSuccessors(s) algorithm, where line $11-14$ (the highlighted part) differs from the original algorithm. In SIPP/C, the collision intervals, the complements of safety intervals, are added to vertices and edges as follows:
SIPP/C vertex conflict: for a robot at the vertex $u_{k}$ in a planned path at time step $k$, we add the collision interval $[k,k]$ to vertices and edges that intersect with the robot-robot collision model at $pos(u_{k})$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$, $\forall v,e$. Note here, we use the swept model $\mathcal{R}^{*}_{\mathcal{R}}(e)$ when the robot traverses an edge.
SIPP/C edge conflict: for a robot traversing an edge $e_{k}\!\!:=\!\!(u_{k},u_{k+1})$ at time step $k$, we add the collision interval $[k,k]$ to vertices and edges that intersect with $\mathcal{R}^{*}_{\mathcal{R}}(e_{k})$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}^{*}_{\mathcal{R}}(e_{k})$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}^{*}_{\mathcal{R}}(e_{k})$, $\forall v,e$.
@@ 16             $s^{{}^{\prime}}$ = state of configuration $cfg$ with interval $i$ and time $t$
In Algo. 2, $E(s)$ is the action space at state $s$. For a discrete path $\mathcal{P}^{i}$, we assign a time $t_{k}=k\Delta t$ to each discrete time step and obtain the path $f^{i}$. $\Delta t$ is a user-defined value to satisfy the robot’s dynamic constraints. The low-level replanning happens repeatedly with a $\delta_{l}$ time interval.
@@ VII-C Cell-crossing Protocol
A robot would idle at a local goal if the path in its next cell is not yet computed. We propose a cell-crossing protocol that results in non-stop execution, even when traversing between cells. We buffer the hyperplane $H_{ml}$ by a distance $d_{e}$ towards cell $P_{m}$. All robots, currently in $P_{m}$ and exiting to $P_{l}$, compute paths for $P_{l}$ within the buffer. Buffering is achieved by changing the hyperplane offset to $\mathcal{H}^{{}^{\prime}}_{a}=\mathcal{H}_{a}-\mathcal{H}_{\mathbf{n}}\cdot d_{e}$. By enforcing the buffer distance $d_{e}\geq\delta_{l}\cdot V_{max}$, where $V_{max}$ is the maximum robot speed, the robot is guaranteed to have a plan in its next cell computed at least once before leaving its current cell. A robot entering the buffer zone will then have a plan to exit its current cell and transition through its next cell. The robot, in the buffer zone, fixes its plan of the current cell to lock the local goal and expected arrival time to its next cell. Thus, when computing the plan for the next cell, the robot’s expected start time and position will be pre-determined and independent of cell planning order. The robot computes a plan for its next cell, then concatenates the fixed plan of its current cell and the plan in its next cell to form a complete transition plan. Fig. 5 depicts our cell-crossing protocol.
@@ VIII Results and Discussion
Fig. 6 depicts typical solutions of the proposed algorithm. We summarize quantitative results in Table I, where MCF refers to MCF/OD and One-shot MCF running in parallel, as described in Sec. VI-A. Note that, as $\left|V\right|$ and $\left|E\right|$ suggest, the number of vertices and edges increases after partitioning as we add local goals and corresponding edges. With a small computational overhead $\bar{t}_{\mathrm{high}}$, the proposed high-level planner effectively reduces the congestion among cells compared to both greedy and partitionless baseline approaches, by inspecting $\bar{N}_{\mathrm{max}}$. Here $\bar{t}_{\mathrm{high}}$ is the average high-level planning time and $\bar{N}_{\mathrm{max}}$ is the averaged maximum number of robots in a cell throughout the whole execution. By increasing the number of cells $Q$, the algorithm significantly reduces MAPF computation time. Specifically, for the average low-level planning time $\bar{t}_{\mathrm{low}}$, in instance “Circle74”, the proposed algorithm runs 616-times faster than the baseline method and 544-times faster in instance “Circle142”. Both the average low-level replanning time $\bar{t}_{\mathrm{low}}$ and the averaged maximum low-level replanning time $\bar{t}^{\mathrm{max}}_{\mathrm{low}}$ are within the real-time regime for all instances. Since we use an anytime algorithm, we only record its initial planning time in Table I, which can be directly compared to the baseline. As we would expect, while running in real-time, the algorithm yields suboptimal solutions compared to the baseline MAPF, according to the average makespan $\bar{T}$. Because the partitioning invalidates the global optimality of MAPF algorithms, and the high-level planner detours robots among the partition, the planned paths become longer.
@@ VIII-A Effectiveness of Partition in Low-level Planning
Fig. 7a shows a quantitative evaluation of the low-level planning time and its solution makespan against the number of cells in “Circle74”. We report the median of low-level planning time for its robustness. The logarithm of low-level planning time decreases significantly as the number of cells increases. Thus, by dividing the workspace and parallelizing computation, the proposed algorithm significantly reduces the low-level planning time. The makespan increases at the beginning then plateaus. This phenomenon is expected. As the number of cells increases, at the beginning, the high-level planner becomes more effective in detouring the robots to reduce congestion. At a certain point, no further detour is required as congestion regulation is satisfied. Additionally, the partition breaks the global optimality of the MAPF planner, leading to a degeneracy in solution quality.
@@
Our algorithm runs MAPF in parallel for all cells; in Fig. 7b, we investigate the utilization of multi-threading. The orange dashed curve represents the theoretical lower bound of the total low-level planning time (assuming no overhead is introduced when allocating the computation to different threads). We observe that our method’s computation time follows the same trend, while the gap between the theoretical lower bound and experimental computation time increases as we increase the number of threads, indicating an increase in parallelization overhead in more cells.